Which number would not be one of the possible solutions to this inequality?8t-12>16t+4.a•-3 b•-4 c•-2 d•-12

Answer by rothauserc(4718)

we have 8t-12>16t+4 combining terms we have
8t>16t+16 or t>2t+2 using this inequality we evaluate it for answers a,b,c,d
a) -3 > -6+2 = -4
b) -4 > -8+2 = -6
c) -2 > -4+2 = -2
d) -12 > -24+2 = -22
note that our inequality is true for a, b, d but NOT for c
-2 is not a possible solution to our inequality